OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

OpenStudy (texaschic101):

-2x + y = 1---->(-1) -4x + y = -1 ------------- 2x - y = -1 (result of multiplying by -1) -4x + y = -1 -------------add -2x = -2 x = 1 now sub 1 in for x in either equation to find y -2x + y = 1 -2(1) + y = 1 -2 + y = 1 y = 1 + 2 y = 3 answer : x = 1 and y = 3......(1,3)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

nope, Thank you :)

OpenStudy (texaschic101):

no problem :)